OpenStudy (anonymous):

Use the quadratic formula to solve the equation. 4x^2-6x-10=0

OpenStudy (anonymous):

factoring out 2 we have 2x^2-3x-5=0 which is (x+1)(2x-5)=0 ang so x=5/2 or x=-1

OpenStudy (anonymous):

it is longer with the quadratic f bu simply substitute the coeficients in x=

OpenStudy (anonymous):

I got it, thank you